# Hölder regularity of Siciak-Zaharjuta extremal functions on compact Hermitian manifolds

### 02 Apr 2026 v1

@@ Abstract.
For a compact subset in a compact Hermitian manifold, we prove that the Hölder continuity of the extremal function at a given point in the set is a local property and that the Hölder continuity of a weighted extremal function follows from the Hölder continuities of the extremal function and the weight function with a uniform density in capacity. The second result can be seen as a continuation of a result of Lu, Phung and Tô [LPT21]. Moreover, for a compact subset in a compact Hermitian manifold, we prove that the Hölder continuity of the extremal function with the uniform density in capacity is equivalent to the local Hölder continuity property. These results are generalizations of the results of Nguyen [Ng24] on compact Kähler manifolds. We also show that the $\mu$-Hölder continuity property of a convex compact subset in $\mathbb{C}^{n}$ implies the local $\mu$-Hölder continuity property of order $1$.
@@ 1. Introduction
The Hölder continuity of the solution to the complex Monge-Ampère equations is the best global regularity when the density is in $L^{p}$ with $p>1$. (Let $\|\cdot\|$ be the standard Euclidean norm on $\mathbb{C}^{n}$. For each $p\in(1,2)$ and $\beta\in(2-2/p,1)$, $u(z):=\|z\|^{\beta}$ has $L^{p}$-density, but $u$ is not Lipschitz continuous at $\mathbf{0}$.) Kołodziej [Ko03] first proved the Hölder continuity with an unspecified Hölder exponent. After that, Demailly, Dinew, Guedj, Pham, Kołodziej and Zeriahi in [DDGHKZ] showed it with an explicit exponent on compact Kähler manifolds. After that, Kołodziej and Nguyen in [KN19] showed the Hölder continuity on compact Hermitian manifolds with strictly positive density. Also, Kołodziej and Nguyen in [KN18] showed it on compact Hermitian manifolds with non-negative density.
@@
The Siciak-Zaharjuta extremal function can serve as such a subsolution when it is Hölder continuous. Lu, Phung and Tô in [LPT21] also showed that for a Hölder continuous weight function on a compact Hermitian manifold, its $\omega$-plurisubharmonic envelope becomes an $\omega$-plurisubharmonic Hölder continuous function. We continue the result of [LPT21] for general compact subsets of compact Hermitian manifolds.
@@
Our main theorems are about the Hölder continuity of extremal functions for non-pluripolar compact sets in $X$. Lu, Phung and Tô in [LPT21] proved that for a given Hölder continuous real-valued function $\phi$ on $X$ with a Hölder exponent $\alpha\in(0,1)$, the envelope $V_{X,\phi}^{*}$ is also Hölder continuous with the same exponent $\alpha$. $V_{X,\phi}$ is continuous by [Ahn26a] since $V_{X}\equiv 0$ and $\phi$ are continuous, thus $V_{X,\phi}^{*}=V_{X,\phi}$. However, the Hölder exponent or even Hölder continuity of the weighted extremal functions for a general non-pluripolar compact subset is not always preserved. We need an additional condition, called the uniform density in capacity [Ng24].
@@
Let $0<\mu,\nu\leq 1$ and $0<q,q_{0}$ be constants. In Theorem 1.2, weak local $\mu$-HCP of order $q$ implies a uniform density in capacity of order $nq$ and $2\mu/(q+2)$-Hölder continuity. Conversely, $\nu$-Hölder continuity and a uniform density in capacity of order $q_{0}$ imply local $\nu$-HCP of order $2q_{0}+2$. Also, $\mu$-HCP of order $q$ implies weak local $\mu$-HCP of order $q$ by definitions. Conversely, in Remark 5.4, we give the proof that weak local $\mu$-HCP of order $q$ directly implies local $\mu$-HCP of order $(n+1)q$.
@@
For $\mathbb{K}=\mathbb{R}$ or $\mathbb{K}=\mathbb{C}$, the closure of a bounded Lipschitz domain in $\mathbb{K}^{n}$ and the closure of a bounded domain in $\mathbb{K}^{n}$ with the geometrical condition are shown to have the local Hölder continuity property as subsets of $\mathbb{C}^{n}$ in [Ng24, Example 4.5]. In particular, the real cube $0,1]^{n}$, the polydisk $\bar{\mathbb{B}}(\mathbf{0},1)^{n}$ and a convex compact subset of $\mathbb{K}^{n}$ with the nonempty interior in $\mathbb{K}^{n}$ has a uniform density in capacity in $\mathbb{C}^{n}$. In [Ng24, Example 4.5], the uniform interior sphere condition implies the geometrical condition since the union of uniform closed balls conjecture was proved by Nour, Stern and Takche [NST11]. [Ng24, Lemma 4.7-(a)] directly showed that the closure of a bounded domain in $\mathbb{C}^{n}$ with a uniform interior sphere condition implies the local Hölder continuity property with improvement in both the exponent and the order compared to the geometric condition. This direct proof is also valid since the uniform interior $r$-sphere condition implies the union of uniform closed $r/2$-balls by [NST11].
@@
Let us go to the research of the Hölder continuity of the extremal functions on $\mathbb{C}^{n}$. The HCP of a general compact subset in $\mathbb{C}^{n}$ (introduced in Section 3) does not imply the local HCP (introduced in Section 3) of the subset . However, if such an compact subset is convex, then the author proves that the $\mu$-HCP of the subset implies the local $\mu$-HCP of it with the order $1$ for each $\mu\in(0,1]$.
@@
Let $F$ be a compact convex subset of $\mathbb{C}^{n}$ and $\mu\in(0,1]$. If $F$ has $\mu$-HCP, then $F$ has local $\mu$-HCP of order $1$.
Nguyen [Ng24, Corollary 4.6] proved that a compact subset $F$ in $\mathbb{K}^{n}$, where $\mathbb{K}$ is either $\mathbb{R}$ or $\mathbb{C}$, with the geometrical condition in $\mathbb{K}^{n}$ has local $1/2$-HCP of order $n$.
@@
Let $F$ be a compact subset of $\mathbb{K}^{n}$, where $\mathbb{K}$ is either $\mathbb{R}$ or $\mathbb{C}$. $F$ is said to satisfy the geometrical condition if there exists a constant $r_{F}>0$ such that, for each $a\in F$, there exists $a^{\prime}\in F$ letting the convex hull of $\{a\}\cup\bar{\mathbb{B}}_{\mathbb{K}^{n}}(a^{\prime},r_{F})$ be contained in $F$.
Such an $F$ in Definition 1.8 automatically has $1/2$-HCP. Then by Theorem 1.7, we get the following corollary.
@@
Let $\mathbb{K}$ be either $\mathbb{R}$ or $\mathbb{C}$. Then a convex compact subset of $\mathbb{K}^{n}$ with the nonempty interior in $\mathbb{K}^{n}$ has local $1/2$-HCP of of order $1$.
@@ 2. Basic properties of extremal functions
The basic properties introduced in this section are the analogues for the extremal functions of the compact sets in $\mathbb{C}^{n}$. For the details of these basic properties, we refer to [Ahn26a]. Note that for subsets $E$ and $O$ of $X$ (resp. of $\mathbb{C}^{n}$), if $O$ is open in $X$ (resp. in $\mathbb{C}^{n}$), then $\sup_{U}V_{E}=\sup_{U}V_{E}^{*}$ (resp. $\sup_{O}L_{E}=\sup_{O}L_{E}^{*}$).
@@ 3.2. Local Hölder continuity property
    Equivalently, $K$ has weakly local $\mu$-HCP of order $q$ if for some finite open cover $\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$ of $X$ of holomorphic coordinate balls of $\tau_{j}(\Omega_{j})=\mathbb{B}(\mathbf{0},5)$ with the refined covering $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}_{j\in\Lambda}$ of $X$, there exists a constant $C^{\prime}>0$ such that, for each $(b,j)\in K\times\Lambda$ of $b\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$,
@@
In Definition 3.1, if $K$ has local $\mu$-HCP of order $q$ at $b\in K$, then letting $\delta\to 0^{+}$ in the inequality (v) implies that $K$ is locally $L$-regular at $b$ in the sense of [Ahn26a]. We in [Ahn26a] showed that local $L$-regularity of $K$ at $b$ gives the continuity of $V_{K}$ at $b$. Thus in this case, we know that $V_{K}$ is continuous at $b$.
Moreover, in Definition 3.1, if $K$ has weak local $\mu$-HCP of order $q$ at $b\in K$, then letting $\delta\to 0^{+}$ in the inequality (v) implies that $K$ is weakly locally $L$-regular at $b$ in the sense of [Ahn26a]. We in [Ahn26a] showed that weak local $L$-regularity of $K$ at $b$ gives the continuity of $V_{K}$ at $b$. Accordingly, even in this weaker condition, we also know the continuity of $V_{K}$ at $b$.
@@
For each $a\in F$ and $0<r\leq 1$, let $T_{a,r\epsilon}:\mathbb{C}^{n}\to\mathbb{C}^{n}$ is the invertible affine transformation given by $T_{a,r\epsilon}(z):=(z-a)/(r\epsilon)+a$. Then we have
@@ 3.4. Relation of extremal functions and uniform density in capacity in the complex coordinate space
Specifically, if (3.4) holds, then (3.5) holds with $q_{1}=q_{0}$ and $A_{1}=\varkappa_{0}^{-1}G(2)$ for some function $G(t)\in(0,\infty)$ only depending on $t\in(0,\infty)$. Conversely, if (3.5) holds, then (3.4) holds with $q_{0}=nq_{1}$ and $\varkappa_{0}=(2\pi/A_{1})^{n}$: for each $(a,r)\in K\times(0,1]$,
@@
Let $K$ be a compact subset of $\mathbb{B}(\mathbf{0},1)\subset\mathbb{C}^{n}$. Suppose $K$ has local $\mu$-HCP of order $q$. Then $K$ has a uniform density in capacity of order $nq$.
@@
We can require $K\subset\mathbb{B}(\mathbf{0},1)$. There exist uniform constants $C>0$ such that
@@ 5.2. Proof of the second theorem
There exists a bi-Lipschitz constant $l>1$ such that $d_{\omega}$ is dominated by $l\times((\tau_{0}\times\tau_{0})^{*}d_{\mathbb{C}^{n}})$ on $\tau_{0}^{-1}(\bar{\mathbb{B}}(\mathbf{0},R_{3}))$. The comparability (6.2), inequality (6.3) and Lemma 6.1 in turn give that, for each $\|z\|\leq\delta\leq R_{3}/l$ and $0<r\leq R_{3}/2$, with
@@
Here, $\delta\in(0,5]$ instead of $\delta\in(0,1]$ is possible due to (3.1). By Remark 3.4, there exist bi-Lipschitz constants $0<l_{1}\leq 1\leq l_{2}$ such that for each $j\in\Lambda$ and $p,p^{\prime}\in\tau_{j}^{-1}(\bar{\mathbb{B}}(\mathbf{0},4))$,
@@
Assume that $V_{K}$ is $\nu$-Hölder continuous and $K$ has a uniform density in capacity for constants $0<\nu\leq 1$ and $q_{0}>0$. We will show that $K$ has the local HCP. Fix finite cover $\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$ of $X$ of holomorphic coordinate balls $\tau_{j}(\Omega_{j})=\mathbb{B}(\mathbf{0},5)$ with its refined cover $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}_{j\in\Lambda}$ of $X$. Let $\varkappa_{0}$ be a constant for this cover for the uniform density in capacity (3.2) of $K$. We use the settings and notations same as those of the proof of (vi) $\Rightarrow$ (iv) in Theorem 1.2: $\bar{B}_{j}(b,r)$, $B_{j}(b,r)$, $b_{j}$, $0<l_{1}\leq 1\leq l_{2}$, $d,D>0$, $\rho_{1,p,j}(\cdot):=d\|\tau_{j}(\cdot)-\tau_{j}(p)\|^{2}$, $\rho_{2,p,j}(\cdot):=D\|\tau_{j}(\cdot)-\tau_{j}(p)\|^{2}$.
@@
By the inequality (3.5), since $l_{3}r_{3}/l^{3}=r_{3}/l^{2}<2r_{3}/(l^{2}+l)<r_{3}/l$, for some constant $c^{\prime}(r_{3},l):=G(2r_{3}/(l^{2}+l))>0$, for $c:=(C^{\prime}_{1}r_{3}^{\mu_{1}})^{n}c^{\prime}(r_{3},l)/(2\pi)^{n}>0$,
@@
Therefore, if a compact set $K\subset X$ has the weak local $\mu$-HCP of order $q$ at $b\in K$, then $K$ has local $\mu$-HCP of order $(n+1)q$ at $b\in K$. By the same reason, if $K$ has the weak local $\mu$-HCP of order $q$, then $K$ has local $\mu$-HCP of order $(n+1)q$.

### 05 Apr 2026 v2

@@ Abstract.
For a compact subset in a compact Hermitian manifold, we prove that the Hölder continuity of the extremal function at a given point in the set is a local property and that the Hölder continuity of a weighted extremal function follows from the Hölder continuities of the extremal function and the weight function with a uniform density in capacity. The second result can be seen as an extension of a result of Lu, Phung and Tô [LPT21]. Moreover, for a compact subset in a compact Hermitian manifold, we prove that the Hölder continuity of the extremal function with the uniform density in capacity is equivalent to the local Hölder continuity property, which is also equivalent to the weak local Hölder continuity property. These results are generalizations of the results of Nguyen [Ng24] on compact Kähler manifolds. We also show that the $\mu$-Hölder continuity property of a convex compact subset in $\mathbb{C}^{n}$ implies the local $\mu$-Hölder continuity property of order $1$.
@@ 1. Introduction
The Hölder continuity of the solution to the complex Monge-Ampère equations is the best global regularity when the density is in $L^{p}$ with $p>1$. (Let $\|\cdot\|$ be the standard Euclidean norm on $\mathbb{C}^{n}$. For each $p\in(1,2)$ and $\beta\in(2-2/p,1)$, $u(z):=\|z\|^{\beta}$ has $L^{p}$-density, but $u$ is not Lipschitz continuous at $\mathbf{0}$.) Kołodziej [Ko08] first proved the Hölder continuity. After that, Demailly, Dinew, Guedj, Pham, Kołodziej and Zeriahi in [DDGHKZ] showed it with an explicit exponent on compact Kähler manifolds. After that, Kołodziej and Nguyen in [KN19] showed the Hölder continuity on compact Hermitian manifolds with strictly positive density. Also, Kołodziej and Nguyen in [KN18] showed it on compact Hermitian manifolds with non-negative density.
@@
The Siciak-Zaharjuta extremal function can serve as such a subsolution when it is Hölder continuous. Lu, Phung and Tô in [LPT21] also showed that for a Hölder continuous weight function on a compact Hermitian manifold, its $\omega$-plurisubharmonic envelope becomes an $\omega$-plurisubharmonic Hölder continuous function. One of our results extends this result of [LPT21] for general compact subsets of compact Hermitian manifolds.
@@
Our main theorems are about the Hölder continuity of extremal functions for non-pluripolar compact sets in $X$. Lu, Phung and Tô in [LPT21] proved that for a given Hölder continuous real-valued function $\phi$ on $X$ with a Hölder exponent $\alpha\in(0,1)$, the envelope $V_{X,\phi}^{*}$ is also Hölder continuous with the same exponent $\alpha$. $V_{X,\phi}$ is continuous by [Ah26] since $V_{X}\equiv 0$ and $\phi$ are continuous, thus $V_{X,\phi}^{*}=V_{X,\phi}$. However, the Hölder exponent or even Hölder continuity of the weighted extremal functions for a general non-pluripolar compact subset is not always preserved. We need an additional condition, called the uniform density in capacity [Ng24].
@@
Let $0<\mu,\nu\leq 1$ and $0<q,q_{0}$ be constants. In Theorem 1.2, the weak local $\mu$-Hölder continuity property of order $q$ implies the uniform density in capacity of order $nq$ and the $2\mu/(q+2)$-Hölder continuity. Conversely, the $\nu$-Hölder continuity and the uniform density in capacity of order $q_{0}$ imply the local $\nu$-Hölder continuity property of order $2q_{0}+2$. Also, the $\mu$-Hölder continuity property of order $q$ implies the weak local $\mu$-Hölder continuity property of order $q$ by the definitions. Conversely, by applying Theorem 1.2 twice, we know that the weak local $\mu$-Hölder continuity property of order $q$ implies the local $2\mu/(q+2)$-Hölder continuity property of order $2nq+2$. In Remark 5.4, we directly prove that the weak local $\mu$-Hölder continuity property of order $q$ directly implies the local $\mu$-Hölder continuity property of order $(n+1)q$.
@@
For $\mathbb{K}=\mathbb{R}$ or $\mathbb{K}=\mathbb{C}$, the closure of a bounded Lipschitz domain in $\mathbb{K}^{n}$ and the closure of a bounded domain in $\mathbb{K}^{n}$ with the geometrical condition, explained in Definition 1.8, are shown to have the local Hölder continuity property as subsets of $\mathbb{C}^{n}$ in [Ng24, Example 4.5]. In particular, the real cube $0,1]^{n}$, the polydisk $\bar{\mathbb{B}}(\mathbf{0},1)^{n}$ and a convex compact subset of $\mathbb{K}^{n}$ with the nonempty interior in $\mathbb{K}^{n}$ has a uniform density in capacity in $\mathbb{C}^{n}$. In [Ng24, Example 4.5], the uniform interior sphere condition implies the geometrical condition since the union of uniform closed balls conjecture was proved by Nour, Stern and Takche [NST11]. [Ng24, Lemma 4.7-(a)] directly showed that the closure of a bounded domain in $\mathbb{C}^{n}$ with a uniform interior sphere condition implies the local Hölder continuity property with improvement in both the exponent and the order compared to the geometric condition. This direct proof is also valid since the uniform interior $r$-sphere condition implies the union of uniform closed $r/2$-balls by [NST11].
@@
The Hölder continuity property of a general compact subset in $\mathbb{C}^{n}$, introduced in Section 3, does not imply the local Hölder continuity property, introduced in Section 3, of the subset. However, if a compact subset in $\mathbb{C}^{n}$ is convex, then the author proves that the Hölder continuity property of the subset implies the local Hölder continuity property of it with the same exponent and the order $1$.
@@
Let $F$ be a convex compact subset of $\mathbb{C}^{n}$ and $\mu\in(0,1]$. If $F$ has $\mu$-Hölder continuity property, then $F$ has local $\mu$-Hölder continuity property of order $1$.
Nguyen [Ng24, Corollary 4.6] proved that a compact subset $F$ in $\mathbb{K}^{n}$, where $\mathbb{K}$ is either $\mathbb{R}$ or $\mathbb{C}$, with the following geometrical condition in $\mathbb{K}^{n}$ has local $1/2$-Hölder continuity property of order $n$.
@@
Let $F$ be a compact subset of $\mathbb{K}^{n}$, where $\mathbb{K}$ is either $\mathbb{R}$ or $\mathbb{C}$. $F$ is said to satisfy the geometrical condition if there exists a constant $r_{F}>0$ such that, for each $a\in F$, there exists $a^{\prime}\in F$ allowing the convex hull of $\{a\}\cup\bar{\mathbb{B}}_{\mathbb{K}^{n}}(a^{\prime},r_{F})$ to be contained in $F$. Here, ${\bar{\mathbb{B}}}_{\mathbb{K}^{n}}(a^{\prime},r_{F})$ is the closed Euclidean ball in $\mathbb{K}^{n}$ centered at $a^{\prime}$ of radius $r_{F}$.
Then by Theorem 1.7, we get the following corollary.
@@
Let $\mathbb{K}$ be either $\mathbb{R}$ or $\mathbb{C}$. Then a convex compact subset of $\mathbb{K}^{n}$ with nonempty interior in $\mathbb{K}^{n}$ has local $1/2$-Hölder continuity property of order $1$.
@@ 2. Basic properties of extremal functions
The basic properties introduced in this section are the analogues for the extremal functions of the compact sets in $\mathbb{C}^{n}$. For the details of these basic properties, we refer to [Ah26]. Note that for subsets $E$ and $O$ of $X$ (resp. of $\mathbb{C}^{n}$), if $O$ is open in $X$ (resp. in $\mathbb{C}^{n}$), then $\sup_{U}V_{E}=\sup_{U}V_{E}^{*}$ (resp. $\sup_{O}L_{E}=\sup_{O}L_{E}^{*}$).
@@ 3.2. Local Hölder continuity property
    Equivalently, $K$ has weak local $\mu$-HCP of order $q$ if for some finite open cover $\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$ of $X$ of holomorphic coordinate balls of $\tau_{j}(\Omega_{j})=\mathbb{B}(\mathbf{0},5)$ with the refined covering $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}_{j\in\Lambda}$ of $X$, there exists a constant $C^{\prime}>0$ such that, for each $(b,j)\in K\times\Lambda$ of $b\in K\cap\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1))$,
@@
In Definition 3.1, if $K$ has local $\mu$-HCP of order $q$ at $b\in K$, then letting $\delta\to 0^{+}$ in the inequality (v) implies that $K$ is locally $L$-regular at $b$ in the sense of [Ah26]. We in [Ah26] showed that local $L$-regularity of $K$ at $b$ gives the continuity of $V_{K}$ at $b$. Thus in this case, we know that $V_{K}$ is continuous at $b$.
Moreover, in Definition 3.1, if $K$ has weak local $\mu$-HCP of order $q$ at $b\in K$, then letting $\delta\to 0^{+}$ in the inequality (v) implies that $K$ is weakly locally $L$-regular at $b$ in the sense of [Ah26]. We in [Ah26] showed that weak local $L$-regularity of $K$ at $b$ gives the continuity of $V_{K}$ at $b$. Accordingly, even in this weaker condition, we also know the continuity of $V_{K}$ at $b$.
@@
For each $a\in F$ and $0<r\leq 1$, let $T_{a,r\epsilon}:\mathbb{C}^{n}\to\mathbb{C}^{n}$ be the invertible affine transformation given by $T_{a,r\epsilon}(z):=(z-a)/(r\epsilon)+a$. Then we have
@@ 3.4. Relation of extremal functions and uniform density in capacity in the complex coordinate space
Specifically, if (3.4) holds, then (3.5) holds with $q_{1}=q_{0}$ and $A_{1}=\varkappa_{0}^{-1}G(2)$ for some function $G(t)\in(0,\infty)$ only depending on $t\in(0,\infty)$. More generally, by [Ko05, Theorem 2.7], for a compact non-pluripolar subset $F$ of $\mathbb{C}^{n}$, if $0<r<R<\infty$ and $F\subset\mathbb{B}(\mathbf{0},r)$, we have
(3.6) $\frac{2\pi}{Cap(F,\mathbb{B}(\mathbf{0},R))^{1/n}}\leq\sup_{\bar{\mathbb{B}}(\mathbf{0},R)}L_{F}^{*}\leq\frac{G(r)}{Cap(F,\mathbb{B}(\mathbf{0},R))}.$
Conversely, if (3.5) holds, then (3.4) holds with $q_{0}=nq_{1}$ and $\varkappa_{0}=(2\pi/A_{1})^{n}$: for each $(a,r)\in K\times(0,1]$,
@@
Let $K$ be a compact subset of $\mathbb{C}^{n}$. Suppose $K$ has local $\mu$-HCP of order $q$. Then $K$ has a uniform density in capacity of order $nq$.
@@
We can reduce to the case $K\subset\mathbb{B}(\mathbf{0},1)$. There exists a uniform constant $C>0$ such that
@@ 5.2. Proof of the second theorem
There exists a Lipschitz constant $l>1$ such that $d_{\omega}$ is dominated by $l\times((\tau_{0}\times\tau_{0})^{*}d_{\mathbb{C}^{n}})$ on $\tau_{0}^{-1}(\bar{\mathbb{B}}(\mathbf{0},R_{3}))$. The comparability (6.2), inequality (6.3) and Lemma 6.1 in turn give that, for each $\|z\|\leq\delta\leq R_{3}/l$ and $0<r\leq R_{3}/2$, with
@@
Here, $\delta\in(0,5]$ instead of $\delta\in(0,1]$ is possible due to (3.1). By Remark 3.4, there exist Lipschitz constants $0<l_{1}\leq 1\leq l_{2}$ such that for each $j\in\Lambda$ and $p,p^{\prime}\in\tau_{j}^{-1}(\bar{\mathbb{B}}(\mathbf{0},4))$,
@@
Assume that $V_{K}$ is $\nu$-Hölder continuous and $K$ has a uniform density in capacity for constants $0<\nu\leq 1$ and $q_{0}>0$. We will show that $K$ has the local HCP. Fix finite cover $\{(\Omega_{j},\tau_{j})\}_{j\in\Lambda}$ of $X$ of holomorphic coordinate balls $\tau_{j}(\Omega_{j})=\mathbb{B}(\mathbf{0},5)$ with its refined cover $\{(\tau_{j}^{-1}(\mathbb{B}(\mathbf{0},1)),\tau_{j})\}_{j\in\Lambda}$ of $X$. Let $\varkappa_{0}$ be a constant for this cover for the uniform density in capacity (3.2) of $K$. We use the settings and notations same as those of the proof of (vi) $\Rightarrow$ (iv) in Theorem 1.2: $\bar{B}_{j}(b,r)$, $B_{j}(b,r)$, $b_{j}$, $0<l_{1}\leq 1\leq l_{2}$, $d>0$, $D>0$, $\rho_{1,p,j}(\cdot):=d\|\tau_{j}(\cdot)-\tau_{j}(p)\|^{2}$, $\rho_{2,p,j}(\cdot):=D\|\tau_{j}(\cdot)-\tau_{j}(p)\|^{2}$.
@@
By the inequality (3.6), since $lr_{3}/l^{3}=r_{3}/l^{2}<2r_{3}/(l^{2}+l)<r_{3}/l$, for some constant $c^{\prime}(r_{3},l):=G(2r_{3}/(l^{2}+l))>0$, for $c:=(C^{\prime}_{1}r_{3}^{\mu_{1}})^{n}c^{\prime}(r_{3},l)/(2\pi)^{n}>0$,
@@
Therefore, for constants $0<\mu\leq 1$ and $q>0$, if a compact set $K\subset X$ has the weak local $\mu$-HCP of order $q$ at $b\in K$, then $K$ has local $\mu$-HCP of order $(n+1)q$ at $b\in K$. By the same reason, if $K$ has the weak local $\mu$-HCP of order $q$, then $K$ has local $\mu$-HCP of order $(n+1)q$.